OpenStudy (anonymous):

Prove: tan^2xcosx/2secx=1/2sin^2x

OpenStudy (anonymous):

\[(\tan ^2 x)(cosx)\div2secx = (1/2)\sin^2x\]

OpenStudy (xapproachesinfinity):

start by using tanx =sinx/cosx also secx=1/cosx

OpenStudy (xapproachesinfinity):

you should have something like this : \[\frac{\sin^2 x}{\cos^2 x}\frac{\cos x}{2}cos x=?\] this after using the identities i mentioned above

OpenStudy (xapproachesinfinity):

it is all clear from there you should be able to carry it out

OpenStudy (xapproachesinfinity):

just one mention one thing if \[\sec x=\frac{1}{\cos x}\] then \[\cos x=\frac{1}{\sec x}\] just and inverse relation between cos and sec i used this if you read my comment above